OpenStudy (anonymous):

The distance on the number line from -10 to -2 equals 5x – 12. What is the value of x?

OpenStudy (anonymous):

Distance will be: -2 - (-10) = -2 + 10 = 8 So, 5x - 12 = 8 Add 12 to both the sides and tell me what you will get??